A is a 2x2 matrix

Let A = {(2) (0) (0) (-1)}. Find A32

1
Expert's answer
2021-08-04T18:11:44-0400
"A=\begin{pmatrix}
 2 & 0 \\
 0 & -1
\end{pmatrix}"

"A^2=\begin{pmatrix}
 2 & 0 \\
 0 & -1
\end{pmatrix}\begin{pmatrix}
 2 & 0 \\
 0 & -1
\end{pmatrix}=\begin{pmatrix}
 2^2 & 0 \\
 0 & (-1)^2
\end{pmatrix}"

"=\begin{pmatrix}
 4 & 0 \\
 0 & 1
\end{pmatrix}"

"A^4=(A^2)^2=\begin{pmatrix}
 4 & 0 \\
 0 & 1
\end{pmatrix}\begin{pmatrix}
 4 & 0 \\
 0 & 1
\end{pmatrix}=\begin{pmatrix}
 2^4 & 0 \\
 0 & (-1)^4
\end{pmatrix}"

"=\begin{pmatrix}
 16 & 0 \\
 0 & 1
\end{pmatrix}"

"A^8=(A^4)^2=\begin{pmatrix}
 16 & 0 \\
 0 & 1
\end{pmatrix}\begin{pmatrix}
 16 & 0 \\
 0 & 1
\end{pmatrix}=\begin{pmatrix}
 2^8 & 0 \\
 0 & (-1)^8
\end{pmatrix}"

"=\begin{pmatrix}
 256 & 0 \\
 0 & 1
\end{pmatrix}"


"A^{16}=(A^8)^2=\begin{pmatrix}
 256 & 0 \\
 0 & 1
\end{pmatrix}\begin{pmatrix}
 256 & 0 \\
 0 & 1
\end{pmatrix}=\begin{pmatrix}
 2^{16} & 0 \\
 0 & (-1)^{16}
\end{pmatrix}"

"=\begin{pmatrix}
 65536 & 0 \\
 0 & 1
\end{pmatrix}"


"A^{32}=(A^{16})^2=\begin{pmatrix}
 65536 & 0 \\
 0 & 1
\end{pmatrix}\begin{pmatrix}
 65536 & 0 \\
 0 & 1
\end{pmatrix}"

"=\begin{pmatrix}
 2^{32} & 0 \\
 0 & (-1)^{32}
\end{pmatrix}=\begin{pmatrix}
 4294967296 & 0 \\
 0 & 1
\end{pmatrix}"